OpenStudy (anonymous):

A cylindrical can, open at the top, is to hold 500 cm3 of liquid. Find the height and radius that minimize the amount of material needed to manufacture the can.

OpenStudy (ddcamp):

Height = h Radius = r Volume = 500 = h×(pi)r² SA = 2(pi)r² + 2(pi)rh Where I would start: \[h=\frac{ 500}{ \pi r^2 }\]